# Learning parameter dependence for Fourier-based option pricing with tensor networks

### v2

@@ V.2 Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi^{\phi}_{\mathrm{max}}$ and $\chi^{\hat{v}}_{\mathrm{max}}$, respectively, in TABLE 2.
@@ V.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ could generally be maintained at less than 10 (refer to TABLE 2). The details of this compression by SVD are described in Appendix A. Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
We argue that, when the computational complexity of the two approaches is at least comparable, the TT-based option pricing has an advantage in terms of computational time. This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions. Indeed, in our demonstration, we have shown that the TT-based method is advantageous in terms of computational complexity and exhibits even better performance in terms of computational time. However, it is important to note that the actual wall clock time may vary depending on the quality of the implementation.
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $4.76\times 10^{4}$ $5.0\times 10^{5}$ $3.27\times 10^{-4}$ $2.84\times 10^{-2}$ $4$ $7$
$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $7.94\times 10^{4}$ $6.0\times 10^{5}$ $6.11\times 10^{-4}$ $3.38\times 10^{-2}$ $5$ $7$
$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $1.21\times 10^{5}$ $7.0\times 10^{5}$ $7.66\times 10^{-4}$ $3.71\times 10^{-2}$ $6$ $7$
$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $1.48\times 10^{5}$ $8.0\times 10^{5}$ $9.07\times 10^{-4}$ $4.09\times 10^{-2}$ $6$ $7$
$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $2.01\times 10^{5}$ $9.0\times 10^{5}$ $1.24\times 10^{-3}$ $4.51\times 10^{-2}$ $7$ $7$
$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $2.48\times 10^{5}$ $1.0\times 10^{6}$ $2.19\times 10^{-3}$ $5.11\times 10^{-2}$ $7$ $7$
$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $3.02\times 10^{5}$ $1.1\times 10^{6}$ $2.57\times 10^{-3}$ $5.04\times 10^{-2}$ $8$ $7$
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $3.53\times 10^{4}$ $5.0\times 10^{5}$ $2.47\times 10^{-4}$ $2.90\times 10^{-2}$ $4$ $6$
$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $6.38\times 10^{4}$ $6.0\times 10^{5}$ $5.61\times 10^{-4}$ $3.28\times 10^{-2}$ $5$ $6$
$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.00\times 10^{5}$ $7.0\times 10^{5}$ $6.80\times 10^{-4}$ $3.66\times 10^{-2}$ $6$ $7$
$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $1.54\times 10^{5}$ $8.0\times 10^{5}$ $9.71\times 10^{-4}$ $3.93\times 10^{-2}$ $7$ $7$
$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $1.99\times 10^{5}$ $9.0\times 10^{5}$ $1.37\times 10^{-3}$ $4.54\times 10^{-2}$ $7$ $7$
$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $3.00\times 10^{5}$ $1.0\times 10^{6}$ $1.88\times 10^{-3}$ $4.93\times 10^{-2}$ $7$ $8$
$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $7.33\times 10^{5}$ $1.1\times 10^{6}$ $6.25\times 10^{-3}$ $5.28\times 10^{-2}$ $9$ $14$
@@
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{i}\in[0.15,0.25]$ and $S_{0}^{i}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs. (17) and (18), respectively. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
@@
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs. In TT-based option pricing with $\vec{S_{0}}$, the sharp increase in computational complexity at $d=11$ arises. This stems from the larger bond dimensions with a sufficiently low cutoff value for SVD in order to maintain its accuracy
@@
Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$. The odd bond $l=2i-1$ connects the core tensors for $j_{i}$ and $k_{i}$, the indexes for $z_{i}$ and $p_{i}$, respectively, and the even bond $l=2i$ connects those for $k_{i}$ and $j_{i+1}$. In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{i}$ and $z_{i+1}$). It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape. The bond dimensions between sites for $j_{i}$ and $k_{i}$, which are related to the same asset, are large, and the ones between sites for $k_{i}$ and $j_{i+1}$, which are related to different assets, are small.
@@ V.2.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ were reduced to around 10 (refer to Fig. 3(b)). However, at $d$ = 11, the computational complexity of TT-based calculation increases significantly. This is because it was necessary to increase the bond dimension by setting the SVD cutoff sufficiently low to maintain accuracy (refer to TABLE 2). This suggests that $d=11$ may already be reaching the upper limit of $d$ where accuracy can be maintained while surpassing the computational complexity over MC-based calculations.
@@ V.2.4 Total computational time for obtaining the MPOs
Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VI. We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
@@ VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Here, we argue that, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time. This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times. Indeed, in our demonstration, we confirmed that the TT-based method excels in computational complexity and is even more superior in terms of computational time. Note that both TT-based and MC-based option pricing implementations have room for improvement in computational time, and we consider the current comparison in time as reference data. We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use these MPOs to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
@@
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It is a promising approach to directly learning the MPOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.

### v3

@@ V.2 Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{\hat{v}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{K}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
@@ V.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10. The details of this compression by SVD are described in Appendix A. The maximum bond dimension $\chi_{K}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$. Still, the computational complexity is much lower than that of MC-based option pricing. The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $3.12\times 10^{-5}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $3.58\times 10^{-5}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.13\times 10^{-5}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $4.58\times 10^{-5}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.15\times 10^{-5}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0309 0.0357 $2359$ $1.0\times 10^{6}$ $6.68\times 10^{-4}$ $5.11\times 10^{-2}$ $7$ $7$ $19$
$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $2797$ $1.1\times 10^{6}$ $7.91\times 10^{-4}$ $5.04\times 10^{-2}$ $8$ $7$ $20$
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.37\times 10^{-5}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.26\times 10^{-5}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $3.61\times 10^{-5}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.25\times 10^{-5}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $5.12\times 10^{-5}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $6.02\times 10^{-5}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0379 0.0401 $853$ $1.1\times 10^{6}$ $3.43\times 10^{-4}$ $5.28\times 10^{-2}$ $9$ $14$ $11$
@@
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs. (19) and (20), respectively. The computational complexity and time increases when $\eta$ is set by Eq. (19) or (20). $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
@@
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs. In TT-based option pricing with $\vec{\sigma}$ and $\vec{S_{0}}$, the sharp increase in computational complexity at $d=10$ and $d=11$ arises, respectively. This stems from the necessity to set $\eta$ by Eq. (19) or (20), which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{m}$ for $\vec{p}$.
@@
Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$. The odd bond $l=2i-1$ connects the core tensors for $j_{m}$ and $k_{m}$, the indexes for $z_{m}$ and $p_{m}$, respectively, and the even bond $l=2i$ connects those for $k_{m}$ and $j_{m+1}$. In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$). It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape. The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
@@ V.2.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to Fig. 3(b)). The maximum bond dimension, $\chi_{K}$, of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is about 6 when $\eta$ is constant, while it increases to 11 when $\eta$ is adjusted per asset, leading to higher computational complexity at about $d=11$. Nonetheless, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
@@ V.2.4 Total computational time for obtaining the MPOs
Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ and $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VI. We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
@@ VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Here, we argue that, in principle, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time. This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times. However, in our demonstration, compared to the huge computational complexity advantage, the computational time advantage of TT-based option pricing is milder. This is because the TT-based method is bottlenecked by specifying the values of $\vec{p}$. Specifically, accessing the memory for the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ dominates the computational time for additions and multiplications in tensor contraction. Therefore, in the TT-based method, and perhaps the MC-based method too, the implementation may have room for improvement to reduce computational time.
We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the MPO to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
@@
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It is a promising approach to directly learning the MPOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.